# A Weighted Llarull Type Theorem

### v1

@@ Abstract.
This paper generalizes Llarull’s classical scalar curvature rigidity theorem to the setting of weighted manifolds with P-scalar curvature. More precisely, We obtain the P-scalar curvature rigidity of the round sphere. Furthermore, we prove the improvement of Llarull’s theorem due to Listing [15] for P-scalar curvature.
@@ 1. Introduction
([16, Theorem C]) Let $(M,g,)$ be a closed spin Riemannian manifold of dimension $n\geq 3$ with $R(g)\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry.
@@
([15, Theorem 1 ]) Let $(M^{n},g)$ be a closed spin Riemannian manifold of dimension $n\geq 3$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(g)\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry.
Here $\|\wedge^{2}h_{*}\|$ denotes the operator norm of the second exterior power of the differential of $h$, see Definition 1.3. For the proof of this theorem, we also refer to Theorem 3.1 in [20, Theorem 3.1].
The proofs of Theorem 1.1 and Theorem 1.2 all rely on the Dirac operator method, requiring the hypothesis that M is spin. An open question is whether the spin assumption can be removed in both Theorem 1.1 and Theorem 1.2. Recently, Cecchini-Wang-Xie-Zhu [4] proved Theorem 1.1 without the spin assumption in 4-dimension provided $h$ is a distance non-increasing map.
A weighted Riemannian manifold is a Riemannian manifold $(M,g)$ equipped with a smooth function $f:M\to R$ and a measure $e^{-f}dV_{g}$ on $M$. It was first introduced by Lichnerowicz in [13, 14]. Weighted manifolds have appeared for a long time in differential geometry. They play an important role in Perelman’s approach to the Ricci flow. A crucial idea of Perelman’s spectacular proofs [17] required considering manifolds with density and their evolution. This led him to introduce the notion of weighted scalar curvature
$R_{f}(g)\coloneqq R(g)+2\Delta_{g}f-|\nabla f|^{2}_{g}\,$,
@@
As a corollary, we have the following weighted Llarull type theorem.
@@
Let $(M,g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$ with $R_{f}\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry. Consequently, $f$ is a constant.
@@
Excluding Theorem 1.5, one can follow Llarull’s approach by using the weighted Dirac operator method to prove Theorem 1.7 without essential difficulties.
The paper is organized as follows. In section 2, we give some facts about the spin manifolds and weighted Dirac operator. In section 3, we prove Theorem 1.5.
@@ 2. Spin manifold and Weighted Dirac Operator
As standard Dirac operator, the weighted Dirac operator is self-adjoint with respect to the weighted $L^{2}$-inner product. By slight abuse of notation, we denote the connection on $S\otimes E$ by $\nabla$.
@@
The weighted Laplacian $\Delta_{f}$ is a nonpositive operator and is self-adjoint. In fact,
@@
In this section, We adopt Llarull’s setting to prove the main theorem in two cases: even-dimensional case and odd-dimensional case as described in [16].
@@
Next, we follow [20] to prove $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ by contradiction. Assume to the contrary that the inequality $R_{f}\geq(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ is strict somewhere. Then there exists $x_{0}\in M$ and two positive constants $\delta,\delta_{0}>0$ such that for any $y\in U_{\delta}(x_{0})$ with
@@
Since $D_{f,E}\Psi=0$, combining equation (3.3) with equation (3.4), it follows that
@@
It follows from equation (3.7) taht
@@
Plug $u\nabla f=-\frac{B}{A}\nabla u$ and equation (3.10) into equation (3.6) for k=0,m=0k=0,\,m=0, we obtain
@@
Remark 3.1.
For $n\geq 2$, the Theorem 1.5 also holds if one replaces the operator norm $\|\wedge^{2}h_{*}\|$ by $\|h_{*}\|^{2}$.

### v2

@@ Abstract.
This paper generalizes Llarull’s classical scalar curvature rigidity theorem to the setting of weighted manifolds with P-scalar curvature. More precisely, we prove the refinement of Llarull’s theorem for P-scalar curvature, which is similar to Listing’s work [20]. As an application, we establish a Llarull type theorem in the form of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$.
@@ 1. Introduction
([21, Theorem C]) Let $(M,g,)$ be a closed spin Riemannian manifold of dimension $n\geq 3$ with $R(g)\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry.
@@
([20, Theorem 1 ]) Let $(M^{n},g)$ be a closed spin Riemannian manifold of dimension $n\geq 3$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(g)\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry.
Here $\|\wedge^{2}h_{*}\|$ denotes the operator norm of the second exterior power of the differential of $h$, see Definition 1.3. For the proof of this theorem, we also refer to Theorem 3.1 in [25, Theorem 3.1].
The proofs of Theorem 1.1 and Theorem 1.2 all rely on the Dirac operator method, requiring the hypothesis that M is spin. An open question is whether the spin assumption can be removed in both Theorem 1.1 and Theorem 1.2. Recently, Cecchini-Wang-Xie-Zhu [6] proved Theorem 1.1 without the spin assumption in 4-dimension provided $h$ is a distance non-increasing map, Theorem 1.2 plays an important role in their proof.
A weighted Riemannian manifold is a Riemannian manifold $(M,g)$ equipped with a smooth function $f:M\to\mathbb{R}$ and a measure $e^{-f}dV_{g}$ on $M$. It was first introduced by Lichnerowicz in [18, 19]. Weighted manifolds have appeared for a long time in differential geometry. They play an important role in Perelman’s approach to the Ricci flow. A crucial idea of Perelman’s spectacular proofs [22] required considering manifolds with density and their evolution. This led him to introduce the notion of weighted scalar curvature
$R_{f}(g)\coloneqq R(g)+2\Delta_{g}f-|\nabla_{g}f|^{2}_{g}\,$,
@@
Corollary 1.7.
([9, Theorem 1.3]) Let $(M,g,e^{-f}\mathrm{d}V_{g})$ be a weighted closed spin Riemannian manifold of dimension $n\geq 3$ with $R_{f}\geq n(n-1)$. If $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth, area non-increasing map of non-zero degree, then $h$ is an isometry. Consequently, $f$ is a constant.
Remark 1.8.
Deng has already proved Corollary 1.7 for general weighted scalar curvature in [9].
As a corollary of Theorem 1.5, we obtain the Llarull type theorem also holds for warped $\mathbb{T}^{k}$-extension of $M$. see also results of Bär-Brendle-Hanke-Wang in [3] and Wang-Xie in [25] for the case with a non-empty boundary.
@@
Let $(M^{n},g)$ be a spin Riemannian manifold of dimension $n\geq 3$, the warped $\mathbb{T}^{k}$-extension of $M$ is the manifold $(M^{n}\times\mathbb{T}^{k},\bar{g})$ with $\bar{g}=g+\sum_{i=1}^{k}e^{-2f_{i}}d\theta^{2}$, where $f_{i}$ is a smooth function on $M$ for each $1\leq i\leq k$. Assume that $h:(M,g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is a smooth map of non-zero degree such that $R(\bar{g})\geq n(n-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry. Consequently, $f_{i}$ is a constant for each $1\leq i\leq k$.
@@
For $n\geq 2$, Theorem 1.5 and Theorem 1.9 also hold if one replaces the operator norm $\|\wedge^{2}h_{*}\|$ by $\|h_{*}\|^{2}$.
As an application of Theorem 1.9, we generalize Theorem 1.1 in [13] to $\mathbb{S}^{k}$ for $4\leq k\leq 6$ by following the approach of Hao-Shi-Sun in [13].
Theorem 1.11.
Suppose $4\leq n\leq 7$ and $3\leq k\leq n-1$. Let $(M^{n},g)$ be a closed spin Riemannian manifold. If $h:(M,g)\to(\mathbb{S}^{k}\times\mathbb{T}^{n-k},g_{\mathbb{S}^{k}}+g_{\mathbb{T}^{n-k}})$ is a smooth map of non-zero degree such that $R(g)\geq k(k-1)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $(M,g)$ is locally isometric to $(\mathbb{S}^{k}\times\mathbb{T}^{n-k},\frac{1}{c}\cdot(g_{\mathbb{S}^{k}}+g_{\mathbb{T}^{n-k}}))$.
Here, we would like to make a statement: The Theorem 1.11 in the case of $h$ is area non-increasing was also obtained independently by Chow in a recent preprint [7].
The paper is organized as follows. In section 2, we give some facts about the spin manifolds and weighted Dirac operator. In section 3, we prove Theorem 1.5. In section 4, we prove Theorem 1.9 and Theorem 1.11.
@@ 2. Spin manifold and weighted Dirac operator
As standard Dirac operator, the weighted Dirac operator is also self-adjoint with respect to the weighted $L^{2}$-inner product. By slight abuse of notation, we denote the connection on $S\otimes E$ by $\nabla$.
@@
The weighted Laplacian $\Delta_{f}$ is a non-positive operator and is self-adjoint. In fact,
@@
In this section, We adopt Llarull’s setting to prove Theorem 1.5 in two cases: even-dimensional case and odd-dimensional case as described in [21].
@@
Next, we follow Wang-Xie [25] to prove $R_{f}=(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ by contradiction. Assume to the contrary that the inequality $R_{f}\geq(2m-1)(2m-2)\|\wedge^{2}h_{*}\|$ is strict somewhere. Then there exists $x_{0}\in M$ and two positive constants $\delta,\delta_{0}>0$ such that for any $y\in U_{\delta}(x_{0})$ with
@@
Since $D_{f,E}\Psi=0$, combining inequality (3.3) with inequality (3.4), it follows that
@@
It follows from equation (3.7) that
@@
Plugging $u\nabla f=-\frac{B}{A}\nabla u$ and equation (3.10) into equation (3.6) for k=0,m=0k=0,\,m=0, we obtain
@@
4. The Llarull type theorem in the form of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$
In this section, we follow the approach of Hao-Shi-Sun [13] to establish the rigidity result of $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$. We prove Theorem 1.11, specifically addressing the case $\mathbb{S}^{n-1}\times\mathbb{T}^{1}$. For the case $\mathbb{S}^{k}\times\mathbb{T}^{n-k}$, $2\leq n-k\leq n-3$, we use the torical symmetrization technique to achieve dimension reduction as described in [13, Remark 3.3](see also [27, Proposition 2.2]).
Since Theorem 1.9 is crucial for the proof of Theorem 1.11, we first prove it.
Proof of Theorem 1.9.
By direct calculation, we have
(4.1) $R(\bar{g})=R(g)+2\Delta f-|\nabla f|^{2}-\sum_{i=1}^{k}|\nabla f_{i}|^{2},$
where $f=\sum_{i=1}^{k}f_{i}$. Since $R(\bar{g})\geq n(n-1)\|\wedge^{2}h_{*}\|$, $R(g)+2\Delta f-|\nabla f|^{2}\geq n(n-1)\|\wedge^{2}h_{*}\|$. By Theorem 1.5, we obtain there exists a constant $c>0$ such that $h:(M,c\cdot g)\to(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$ is an isometry and $f$ is a constant. Note $\|\wedge^{2}h_{*}\|=c$, by equation (4.1), $f_{i}$ is a constant for each $1\leq i\leq k$. ∎
Recall Theorem 1.11 for the specific case $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$.
Theorem 4.1.
Suppose $4\leq n\leq 7$ . Let $(M^{n},g)$ be a closed spin Riemannian manifold. If $h:(M,g)\to(\mathbb{S}^{n-1}\times\mathbb{S}^{1},g_{\mathbb{S}^{n-1}}+d\theta^{2})$ is a smooth map of non-zero degree such that $R(g)\geq(n-1)(n-2)\|\wedge^{2}h_{*}\|$, then there exists a constant $c>0$ such that $(M,g)$ is locally isometric to $(\mathbb{S}^{n-1}\times\mathbb{S}^{1},\frac{1}{c}\cdot(g_{\mathbb{S}^{n-1}}+d\theta^{2}))$.
Proof.
Since $h$ has non-zero degree, according to Sard’s theorem and the geometric measure theory, there exist a $\theta_{0}\in\mathbb{S}^{1}$ such that $0\neq[h^{-1}(\mathbb{S}^{n-1}\times\{\theta_{0}\})]\in H_{n-1}(M,\mathbb{Z})$ and a closed 2-sided area-minimizing hypersurface $\Sigma^{n-1}\in[h^{-1}(\mathbb{S}^{n-1}\times\{\theta_{0}\})]$, and $h_{1}\coloneqq P_{1}\circ h|_{\Sigma}:\Sigma\to\mathbb{S}^{n-1}$ has non-zero degree, where $P_{1}:\mathbb{S}^{n-1}\times\mathbb{S}^{1}\to\mathbb{S}^{n-1}$ is the canonical projection. By the second variation of stable minimal hypersurface, the associated Jacobi operator
$J_{\Sigma}\coloneqq-\Delta_{\Sigma}-(\mathrm{Ric}_{g}(\nu,\nu)+|A_{\Sigma}|^{2})=-\Delta_{\Sigma}+\frac{1}{2}(R_{\Sigma}-R(g)-|A_{\Sigma}|^{2})$
is non-negative, where $A_{\Sigma}$ is the second fundamental form of $\Sigma$ and $\nu$ is the unit normal vector of $\Sigma$. Let $\phi$ be the first eigenfunction such that
(4.2) $\begin{split}J_{\Sigma}\phi=&-\Delta_{\Sigma}\phi-(\mathrm{Ric}_{g}(\nu,\nu)+|A_{\Sigma}|^{2})\phi\\ =&-\Delta_{\Sigma}\phi+\frac{1}{2}(R_{\Sigma}-R(g)-|A_{\Sigma}|^{2})\phi\\ =&\lambda\phi,\end{split}$
where $\lambda\geq 0$ is the first eigenvalue. By the non-vanishing Theorem in [12, page 47], $\phi>0$. Since $\Sigma$ has trivial normal bundle, it is also a spin manifold. Now, we consider the manifold $(\Sigma\times\mathbb{S}^{1},g_{\phi}=g|_{\Sigma}+\phi^{2}d\theta^{2})$, which is spin, by direct calculation, we have on $\Sigma$
$R(g_{\phi})=$ $R_{\Sigma}-\frac{2\Delta_{\Sigma}\phi}{\phi}$
$=$ $R(g)+|A_{\Sigma}|^{2}+2\lambda$
$\geq$ $(n-1)(n-2)\|\wedge^{2}(h|_{\Sigma})_{*}\|$
$\geq$ $(n-1)(n-2)\|\wedge^{2}h_{1*}\|.$
Since $h_{1}:\Sigma\to\mathbb{S}^{n-1}$ has non-zero degree, by Theorem 1.9, we deduce that $(\Sigma,g|_{\Sigma})$ is isometric to $(\mathbb{S}^{n-1},\frac{1}{c}\cdot g_{\mathbb{S}^{n-1}})$ for a constant $c>0$, $\|\wedge^{2}h_{1*}\|=c$ and $\phi$ is a constant. Utilizing equation (4.2), we obtain $\lambda=0$, $A_{\Sigma}=0$ and $\mathrm{Ric}_{g}(\nu,\nu)=0$. According to Proposition 3.2 in [4](see also Lemma 3.3 in [27]), we can construct a local foliation $\{\Sigma_{t}\}_{-\epsilon\leq t\leq\epsilon}$ in $M$ such that $\Sigma_{t}$ are of constant mean curvature and $\Sigma_{0}=\Sigma$, let $H_{t}$ denote the mean curvature of $\Sigma_{t}$. We claim that $H_{t}\equiv 0$ for any $t\in(-\epsilon,\epsilon)$. It suffices to prove this for nonnegative $t$, as the argument for the negative side is analogous. For contradiction, suppose that the assertion is not true, then there exists a $t_{0}>0$ such that $H_{t_{0}}=\epsilon_{0}>0$, since otherwise $H_{t}\leq 0$ and the area-minimizing property of $\Sigma$ will imply $H_{t}\equiv 0$. Let $\Omega_{t_{0}}$ denote the region enclosed by $\Sigma$ and $\Sigma_{t_{0}}$ i.e.,
$\Omega_{t_{0}}:=\cup_{t\in[0,t_{0}]}\Sigma_{t}.$
Consider the following functional
(4.3) $E(\Omega)=\int_{\partial^{\ast}\Omega\setminus\Sigma}\mathrm{d}\mathcal{H}^{n-1}-\int_{\Omega_{t_{0}}}\epsilon_{1}(\chi_{\Omega}-\chi_{\bar{\Omega}})\mathrm{d}\mathcal{H}^{n},\,0<\epsilon_{1}<\epsilon_{0},$
for $\Omega,\bar{\Omega}\in\mathcal{C}$ and $\Sigma\subset\bar{\Omega}$, where
$\mathcal{C}=\{\Omega|\mbox{ all Caccioppoli sets }\Omega\subset\Omega_{t_{0}}\mbox{ and }\Omega\triangle\bar{\Omega}\Subset\overset{\circ}{\Omega}_{t_{0}}\}.$
Due to $0<\epsilon_{1}<\epsilon_{0}$, we can find a Borel set $\hat{\Omega}$ minimizing the brane functional $E$ such that $\hat{\Sigma}=\partial\hat{\Omega}\setminus\Sigma$ is a smooth 2-sided hypersurface disjoint from $\Sigma$ and $\Sigma_{t_{0}}$. Let $\hat{\nu}$ and $\hat{A}$ denote the outwards pointing unit normal vector and the second fundamental form of $\hat{\Sigma}$, respectively. Since $\hat{\Sigma}$ is $E$-minimizing, it has constant mean curvature $\epsilon_{1}$ and it is $E$-stable. So the Jacobi operator
$\hat{J}\coloneqq-\Delta_{\hat{g}}-(\mathrm{Ric}_{g}(\hat{\nu},\hat{\nu})+|\hat{A}|^{2})=-\Delta_{\hat{g}}+\frac{1}{2}(R_{\hat{\Sigma}}-R(g)-|\hat{A}|^{2}-\epsilon_{1}^{2})$
is non-negative, where $\hat{g}$ is the induced metric of $\hat{\Sigma}$ from $(M,g)$. Taking the first eigenfunction $\hat{\phi}$ of $\hat{J}$ with respect to the first eigenvalue $\hat{\lambda}\geq 0$ and defining the metric $\tilde{g}=\hat{g}+\hat{\phi}^{2}d\theta^{2}$ on $\hat{\Sigma}\times\mathbb{S}^{1}$, then
$R(\tilde{g})=R(g)+|\hat{A}|^{2}+2\hat{\lambda}+\epsilon_{1}^{2}\geq(n-1)(n-2)\|\wedge^{2}\hat{h}_{1*}\|+\epsilon_{1}^{2},$
where $\hat{h}_{1}\coloneqq P_{1}\circ h|_{\hat{\Sigma}}:\hat{\Sigma}\to\mathbb{S}^{n-1}$, which has non-zero degree since $\hat{\Sigma}$ is homologous to $\Sigma$. By Theorem 1.9, we have $|\hat{A}|=0$ and $\epsilon_{1}=0$, which leads to a contradiction.
Since $H_{t}$ vanishes for any $t\in(-\epsilon,\epsilon),$ there holds $\operatorname{Area}(\Sigma_{t})=\operatorname{Area}(\Sigma)$, which yields that $\Sigma_{t}$ is also area-minimizing. Replacing $\Sigma$ by $\Sigma_{t}$ and combining with Theorem 1.9, $\Sigma_{t}$ is totally geodesic and has vanished normal Ricci curvature.
In the following, we establish the existence of a local isometry $\varPhi:\Sigma\times\mathbb{R}\to M$. Let $\tilde{\nu_{t}}$ and $\tilde{V_{t}}$ denote the outwards pointing unit normal vector field and normal variation vector field of the foliation $\{\Sigma_{t}\}_{-\epsilon\leq t\leq\epsilon}$, respectively. Assume $\varPhi:\Sigma\times(-\epsilon,\epsilon)\to M$ is the flow generated by $\tilde{V_{t}}$. It follows that $\varPhi$ is an embedding in a small neighborhood of $\Sigma$ and the pull-back of the metric $g$ is
$\varPhi^{\ast}(g)=\psi^{2}dt^{2}+\varPhi^{\ast}_{t}(g_{t}),$
where $\psi=\langle\tilde{V_{t}},\tilde{\nu_{t}}\rangle>0$ is the lapse function (see definition in (3.3) in [4]) and $g_{t}$ is the induced metric on $\Sigma_{t}$ from $g$. Since $\Sigma_{t}$ is totally geodesic, we have $\partial_{t}\varPhi^{\ast}_{t}(g_{t})=2\psi A_{\Sigma_{t}}=0$, which implies that $\varPhi^{\ast}_{t}(g_{t})=g|_{\Sigma}$. Additionally, the stability condition yields $\Delta_{\Sigma_{t}}\psi-|A_{\Sigma_{t}}|^{2}\psi-\operatorname{Ric}(\nu_{t},\nu_{t})\psi=-\Delta_{\Sigma_{t}}\psi=0,$ then $\psi(\cdot,t)\equiv\mathrm{constant}$. Let
$r(\cdot,t)=\int_{0}^{t}\psi(\cdot,s)ds,$
then
$\varPhi^{\ast}(g)=dr^{2}+g|_{\Sigma}.$
Hence, $\varPhi:\Sigma\times(-\epsilon,\epsilon)\to M$ is a local isometry. Through a continuous argument as Proposition 3.8 in [4], we conclude that there exists a local isometry $\varPhi:\Sigma\times\mathbb{R}\to M$. Since $(\Sigma,g|_{\Sigma})$ is isometric to $(\mathbb{S}^{n-1},\frac{1}{c}\cdot g_{\mathbb{S}^{n-1}})$ for a constant $c>0$, by scaling the fact $\mathbb{R}$, $(M,g)$ is isometrically covered by $(\mathbb{S}^{n-1}\times\mathbb{R},\frac{1}{c}\cdot(g_{\mathbb{S}^{n-1}}+d\theta^{2}))$. ∎